If the pH of a substance is 4, what will the pOH be?


1
Expert's answer
2022-05-12T03:33:04-0400

Solution:

For any aqueous solution at 25C:

pH + pOH = 14

pOH = 14 - pH = 14 - 4 = 10

pOH = 10


Answer: pOH = 10